# Calibrated and uncertain? Evaluating uncertainty estimates in binary classification models

### v1

@@ Appendix C Identically distributed classes
Figure 12: Estimated probabilities (top row), the absolute difference between the LRFD and the estimated probabilities (middle row) and uncertainties (bottom row) for $N_{\text{train}}=500$ and $N_{\text{train}}=5000$ (bottom) for the different algorithms in the study for dataset A.
@@
Figure 12 shows the spatial distribution of probability estimates, difference between the estimated distributions and the LRFD and the estimated uncertainties. These visualizations nicely illustrate the impact of the inherent spatial structures of the different models. The plain ReLU neural network models like NNE and EDL produce distributions with linear structure. MCD blurs the linear structure, creating a smoother distribution. GP with RBF kernel produces relatively smooth distributions with local radial symmetry. The specific DPMM kernels used in this study are axes-parallel bivariate Gaussians, which give the distinct grid-like structure seen in the plots. These results highlight the importance of choosing appropriate model architecture (or transforming the data to fit the model), even when the modeling tools are in theory very flexible.

### v2

@@ Appendix D Reliability diagrams
Figure 16: Estimated probabilities (top row), the absolute difference between the LRFD and the estimated probabilities (middle row) and uncertainties (bottom row) for $N_{\text{train}}=5000$ using the ReLU activation function (top) and the tanh activation function (bottom) for the deep learning algorithms in the study for dataset A.
@@
Figure 16 shows the spatial distribution of probability estimates, difference between the estimated distributions and the LRFD and the estimated uncertainties and illustrates the impact of changing the activation function in the deep learning architectures. These visualizations nicely illustrate the impact of the inherent spatial structures of the different models. The ReLU neural network models like NNE and EDL produce distributions with linear structure. MCD blurs the linear structure, creating a smoother distribution. The linear structure is even more prominent when using the tanh activation function. GP with RBF kernel produces relatively smooth distributions with local radial symmetry. The specific DPMM kernels used in this study are axes-parallel bivariate Gaussians, which give the distinct grid-like structure seen in the plots. These results highlight the importance of choosing appropriate model architecture (or transforming the data to fit the model), even when the modeling tools are in theory very flexible.